# Bayesian inferences on covariant density functionals from multimessenger astrophysical data: Nucleonic models

### v2

@@ Abstract
We apply the Bayesian inference framework based on CDFs with density dependent couplings. The astrophysical data is expanded to include not only the latest multi-messenger constraints from NICER and gravitational wave events but also the highest measured mass to date for the “black widow” pulsar PSR J0952-0607 and the mass-radius estimates for the ultra-compact, low-mass object HESS J1731-347. The significance of the individual impact of these two sources, as well as the broader picture that emerges from their combined influence, is examined by analyzing the posterior distributions of key CS properties, nuclear characteristics at saturation density, and their correlations.
@@ I Introduction
The systematic construction of dense matter equation of state (EoS) models and statistical inference of EoS parameters from data is an endeavor that has attracted recently significant attention. The Bayesian framework has been applied to models covering the range from fully physics agnostic non-parametric models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] to microscopic models based on nuclear potentials [12, 13, 14, 15, 16] to density functional method-based models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. This effort has matured in the current era of multi-messenger astronomy with a large push to explore model-dependence inferences. The future thus lies in combining more and more sets of data of both types to understand nuclear and neutron star models better.
The main motivation of this work is to extend our previous study of covariant density functionals (CDFs, with linear meson-baryon couplings and density-dependent coupling constants) based Bayesian inference for compact stars (CSs) with well-established astrophysical constraints [31] to those that including more extreme and less certain constraints. This provides a more complete exploration of the parameter space that admits the existence of very massive or ultra-compact stars. The astrophysical constraints for the sources that we considered in the present analysis are summarized in Fig. 1. We aim to assess the compatibility of the parameter space over which these models are defined with the recent multi-messenger observations of CSs. We stress that previous studies of this sort of CDF used simplified functions for density dependence of meson-baryon couplings [18, 19, 24, 28], which thus reduced the flexibility of the model.
The remainder of this introduction provides a detailed overview of current astrophysical constrains in Subsec. I.1 and a review of the current state of the art of the CDFs in Subsec. I.2. Readers familiar with these topics can proceed directly to Sec. II which discusses the specific CDF model of dense matter used in our analysis. The Bayesian inference framework adopted in this work and underlying numerical methods are discussed in Sec. III. Our analysis is specified to the case of ultra-compact low-mass star HESS J1731-347 and the high-mass star PSR J0952-0607 in Sec. IV where we study the impact of these data on the overall picture that emerges once these are included in the inference analysis. Our conclusions and outlook are given in Sec. V.
@@ I.1 Summary of astrophysical constraints
Significant advancements in astrophysical observations of CSs over the past decade have provided valuable constraints on dense matter theories. The first observation of gravitational wave (GWs) event GW170817 from the inspiral of a binary neutron stars by the LIGO-Virgo Collaboration has placed constraints on the dimensionless tidal deformability (TD) of this system, $\tilde{\Lambda}_{\rm GW170817}\leqslant 720$ at 90% confidence interval (hereafter CI), and on the range of the component masses $1.17$-$1.60\,M_{\odot}$ [33, 34, 35]. GW190425 is a highly significant gravitational wave signal, most likely originating from the merger of two neutron stars, making it the second such event ever observed. The TD parameter is constrained to $\tilde{\Lambda}_{\mathrm{GW}190425}\leqslant 600$, with component masses being within the range $1.46$-$1.87\,M_{\odot}$ [36]. Neutron Star Interior Composition Explorer Mission (NICER) has provided mass–radius estimates for four pulsars by observing X-ray emission from hot spots on the neutron stars’ surfaces. The first estimates for one of the most massive known neutron star PSR J0740+6620 [37, 38] was updated from a joint NICER and XMM-Newton analysis of the 2018-2022 dataset, refining the radius and mass to $12.49^{+1.28}_{-0.88}$ km and $2.073^{+0.069}_{-0.069}\,M_{\odot}$, respectively [39]. The updated analysis of the isolated pulsar PSR J0030+0451 revealed that it is a complex source to model, with solutions varying depending on the assumed model. Reference [40] adopted the same models used in the initial studies [41, 42], and found compatible but slightly more stringent results $R=13.71^{+1.30}_{-0.30}$ km and $M=1.37^{+0.17}_{-0.17}\,M_{\odot}$. When XMM-Newton data is accounted for, the model that is more consistent with the magnetic field geometry inferred for the gamma-ray emission for this source, has inferred radius and mass of $R=11.71^{+0.88}_{-0.83}$ km and $M=1.40^{+0.13}_{-0.12}\,M_{\odot}$ (at 68.3% CI). The most complex model tested in Ref. [40], which is highly preferred by the Bayesian evidence, predicts $R=14.438^{+0.883}_{-1.050}$ km and $M=1.70^{+0.18}_{-0.19}\,M_{\odot}$. Recently, new NICER astrophysical constraints have been released for two pulsars — one canonical-mass $1.4\,M_{\odot}$ star PSR J0437-4715 [43], and the one-solar-mass star PSR J1231-1411 [44]. For the nearest and brightest pulsar PSR J0437-4715, its first mass and radius are estimated as $M=1.418^{+0.037}_{-0.037}\,M_{\odot}$ and $R=11.36^{+0.95}_{-0.63}$ km, by using the 2017-2021 NICER X-ray spectral-timing data [43]. The inference results for PSR J1231-1411 exhibit a strong sensitivity to the choice of radius priors, with stable and well-converged outcomes achieved only under constrained radius priors. When the radius is restricted to agree with previous observational constraints and EoS analyses, the inferred values are $R=12.60_{-0.33}^{+0.31}$ km and $M=1.04_{-0.03}^{+0.05}M_{\odot}$ [44].
@@
The mass range for neutron stars covered by the aforementioned data spans the range 1 to $2\,M_{\odot}$. Correspondingly, the relevant range for densities spans from $2$-$4\,\rho_{\rm sat}$, with $\rho_{\rm sat}\simeq 0.16$ fm-3 being the nuclear saturation density. This opens an unprecedented opportunity to explore models of modern theories of dense matter subject to astrophysical constraints. For example, the impact of variations in the estimates for PSR J0030+0451 [40] on dense matter theories and CS properties have recently been assessed [10, 45, 31]. In addition, the X-ray observations of the temperature of the neutron star in the Cassiopeia A supernova remnant and modeling of its thermal evolution give insights into the possible superfluid properties of the core matter [46].
@@
Determining the maximum mass of neutron stars is crucial for understanding stellar evolution, supernovae, and binary mergers, as well as their observational signatures. This value primarily depends on the EoS at high densities, $\rho>4\rho_{\rm{sat}}$. Most observed neutron stars rotate significantly more slowly than their theoretical Keplerian break-up limit (e.g. Ref [47] and references therein). As a result, mass measurements are commonly used to establish lower bounds on the maximum mass of cold, non-rotating neutron stars $M_{\rm max}$, even though rapid rotation can support up to $\sim 20\%$ additional mass. The highest precisely and reliably measured massive mass belongs to pulsar PSR J0348+0432, with a mass estimate of $M=2.01_{-0.04}^{+0.04}\,M_{\odot}$ (at 68.3 % Cl) [48]. The heavier pulsar, PSR J0740+6620, has a mass of $M=2.08_{-0.07}^{+0.07}\,M_{\odot}$ [49, 50] and excludes all the EoS models which cannot reach this limit. At the same time the analysis of the kilonova AT2017gfo and GRB170817A [51, 52], associated with GW170817, have helped in designing theoretical models of post-merger phase of GW170817 that placed an upper limit on the maximum mass $M_{\rm max}\sim 2.3\,M_{\odot}$ [53, 54, 55, 56, 57, 58].
@@
These mass limits have been further challenged by evidence suggesting the existence of even heavier compact objects. The “black widow” pulsar PSR J0952-0607, the second-fastest known spinning CS with a rotation frequency of 707 Hz, has been detected in the Milky Way with a mass estimate of $M=2.35_{-0.17}^{+0.17}\,M_{\odot}$ (at 68.3% CI) [59], making it the most massive neutron star observed to date.
@@ I.2 Summary of nucleonic models of dense matter
The EoS for dense matter in CSs can be constructed using various approaches, including ab initio calculations [73, 74, 75, 76], density functional methods [77, 78, 79, 80], and meta-type models [81, 82, 83], which provide a more schematic representation by parameterizing nuclear saturation properties. These models can also be linked to heavy-ion collision data [84, 85].
@@
The covariant density functional (CDF) models provide a rigorous framework to address the full range of available data on nuclear systems, ranging from the atomic chart to the astrophysics of CSs, for reviews see Refs. [87, 88, 77, 79, 80]. These models were instrumental in addressing successfully such astrophysics problems as hyperon puzzle in conjunction with two-solar mass CSs and TD inference of GW170817 event. The CDF models are separated into broad classes which (a) have non-linear meson contributions to the effective Lagrangian and (b) keep only linear coupling but impose density-dependence of the coupling, which captures the medium modifications of the meson-nucleon vertices. The mesons effectively considered and their self- or cross-couplings in the former class or the explicit forms of density-dependences of the couplings in the latter class are commonly flexible in order to address specific issues. Consequently, their predictions, however, are somewhat model-dependent, as different density functionals result in different dependences on density, see reviews [77, 79, 80].
@@ II Density-dependent covariant density functionals for nucleonic matter
where $\psi_{N}$ is the nucleonic Dirac fields with masses $m_{N}$, $\bm{\tau}$ is the vector of isospin Pauli matrices with $\tau_{3}$ being its third component, and $\sigma,\,\omega_{\mu}$, and $\bm{\rho}_{\mu}$ represent the mesonic fields that mediate the interaction among nucleon fields with the corresponding coupling constants $g_{\sigma},\,g_{\omega}$ and $g_{\rho}$, respectively. The remaining pieces of the Lagrangian correspond to the mesonic, and leptonic contributions, respectively [77, 80].
The particle masses adopted in the present analysis are listed in Table 1, according to the widely used DD-ME parametrization [90, 91]. Note that the mass of $\sigma$ meson $m_{\sigma}$, which is supposed to represent the two-$\pi$-exchange contribution to the nuclear force is not known with precision and lies around 500 MeV. The properties of the infinite nuclear matter described by the present model depend only on the ratios of coupling constants and the corresponding meson masses [92, 93], and our results do not change with variation of $\sigma$ meson mass $m_{\sigma}$, as its variations are equivalently encoded into the variations of $g_{\sigma}$ as fixed $m_{\sigma}$.
Table 1: Nucleon and meson masses (in units of MeV) adopted in the CDF parameterization.
@@
The dependence of meson-nucleon couplings on the total baryon (vector) density is assumed as
@@
which is important for a quantitative description of finite nuclei, e.g., spin-orbit splitting [95]. The rearrangement term $\Sigma_{R}$ in Eqs. (6) arises from the density dependence of meson-baryon couplings and is included to ensure thermodynamical consistency. It seen that at zero temperature, the energy density and pressure of stellar matter are functions of baryonic density only.
It is now clear that if we fix in the Lagrangian (II) the nucleon and meson masses to be (or close to) the ones in the vacuum then properties of infinite nuclear matter can be computed uniquely in terms of seven adjustable parameters. These are the three coupling constants at saturation density $(g_{\sigma},\,g_{\omega},\,g_{\rho})$, and four parameters $(a_{\sigma},\,d_{\sigma},\,d_{\omega},\,a_{\rho})$ that control their density dependence. We further require conditions of weak equilibrium and charge neutrality that prevail in CSs. Combined with the field equations for baryons and mesons, they allow one to determine the equilibrium composition $\rho_{N}$ and $\rho_{l}$ at a given baryon number density and, thus, determine the EoS of stellar matter.
The energy density of isospin asymmetric matter is customarily split into an isoscalar and an isovector term:
@@
where $\chi=(\rho-\rho_{\rm{sat}})/3\rho_{\rm{sat}}$. The coefficients of the expansion are known as nuclear matter characteristics at saturation density, namely, binding energy $E_{\rm{sat}}$, incompressibility $K_{\rm{sat}}$, the skewness $Q_{\rm{sat}}$ in the isoscalar sector, and the symmetry energy $J_{\rm{sym}}$ and its slope of symmetry energy parameter $L_{\rm{sym}}$ in the isovector sector. Given the five macroscopic characteristics together with the preassigned values of $\rho_{\rm sat}$ and $M^{\ast}_{\rm D}$, one can determine uniquely the seven adjustable parameters of the density functional defined above.

### v3

@@ Abstract
We apply the Bayesian inference framework based on CDFs with density dependent couplings. The astrophysical data are expanded to include not only the latest multi-messenger constraints from NICER telescope and gravitational wave events but also the highest measured mass to date for the “black widow” pulsar PSR J0952-0607 and the mass-radius estimates for the ultra-compact, low-mass object HESS J1731-347. The significance of the individual impact of these two sources, as well as the broader picture that emerges from their combined influence, is examined by analyzing the posterior distributions of key CS properties, nuclear characteristics at saturation density, and their correlations.
@@ I Introduction
The systematic construction of dense matter equation of state (EoS) models and statistical inference of EoS parameters from data is an endeavor that has recently attracted significant attention. The Bayesian framework has been applied to models covering the range from fully physics agnostic non-parametric models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] to microscopic models based on nuclear potentials [12, 13, 14, 15, 16] to density functional method-based models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. This effort has matured in the current era of multi-messenger astronomy with a large push to explore model-dependence in statistical inferences. The future thus lies in combining more and more sets of data of both types to understand nuclear and neutron star models better.
The main motivation of this work is to extend our previous study of covariant density functional (CDF, with linear meson-baryon couplings and density-dependent coupling constants) based Bayesian inference for compact stars (CSs) with well-established astrophysical constraints [31] to those that including more extreme and less certain constraints. This provides a more complete exploration of the parameter space that admits the existence of very massive or ultra-compact stars. The astrophysical constraints for the sources that we considered in the present analysis are summarized in Fig. 1. We aim to assess the compatibility of the parameter space over which these models are defined with the recent multi-messenger observations of CSs. We stress that previous studies of this sort of CDF used simplified functions for density dependence of meson-baryon couplings [18, 19, 24, 28], which thus reduced the flexibility of the model.
The remainder of this introduction provides a detailed overview of current astrophysical constrains in Subsec. I.1 and a review of the current state of the art of the dense matter theories in Subsec. I.2. Readers familiar with these topics can proceed directly to Sec. II which discusses the specific CDF model of dense matter used in our analysis. The Bayesian inference framework adopted in this work and underlying numerical methods are discussed in Sec. III. Our analysis is specific to the cases of the ultra-compact low-mass star HESS J1731-347 and the high-mass star PSR J0952-0607 in Sec. IV where we study the impact of these data on the overall picture that emerges once these are included in the inference analysis. Our conclusions and outlook are given in Sec. V.
@@ I.1 Summary of astrophysical constraints
Significant advancements in astrophysical observations of CSs over the past decade have provided valuable constraints on dense matter theories. The first observation of gravitational wave (GW) event GW170817 from the inspiral of a binary neutron stars by the LIGO-Virgo Collaboration placed constraints on the dimensionless tidal deformability (TD) of this system, $\tilde{\Lambda}_{\rm GW170817}\leqslant 720$ at 90% confidence interval (hereafter CI), and on the range of the component masses $1.17$-$1.60\,M_{\odot}$ [33, 34, 35]. GW190425 is a highly significant gravitational wave signal, most likely originating from the merger of two neutron stars, making it the second such event ever observed. The TD parameter is constrained to $\tilde{\Lambda}_{\mathrm{GW}190425}\leqslant 600$, with component masses being within the range $1.46$-$1.87\,M_{\odot}$ [36]. Neutron Star Interior Composition Explorer Mission (NICER) has provided mass–radius estimates for four pulsars by observing X-ray emission from hot spots on the neutron stars’ surfaces. The first estimates for one of the most massive known neutron star PSR J0740+6620 [37, 38] was updated from a joint NICER and XMM-Newton analysis of the 2018-2022 dataset, refining the radius and mass to $12.49^{+1.28}_{-0.88}$ km and $2.073^{+0.069}_{-0.069}\,M_{\odot}$, respectively [39]. The updated analysis of the isolated pulsar PSR J0030+0451 revealed that it is a complex source to model, with solutions varying depending on the assumed model. Reference [40] adopted the same models used in the initial studies [41, 42], and found compatible but slightly more stringent results $R=13.11^{+1.30}_{-1.30}$ km and $M=1.37^{+0.17}_{-0.17}\,M_{\odot}$. When XMM-Newton data are accounted for, the model that is more consistent with the magnetic field geometry inferred for the gamma-ray emission for this source inferred radius and mass of $R=11.71^{+0.88}_{-0.83}$ km and $M=1.40^{+0.13}_{-0.12}\,M_{\odot}$ (at 68.3% CI). The most complex model tested in Ref. [40], which is highly preferred by the Bayesian evidence, predicts $R=14.438^{+0.883}_{-1.050}$ km and $M=1.70^{+0.18}_{-0.19}\,M_{\odot}$. Recently, new NICER astrophysical constraints have been released for two pulsars — one canonical-mass $1.4\,M_{\odot}$ star PSR J0437-4715 [43], and the one-solar-mass star PSR J1231-1411 [44]. For the nearest and brightest pulsar PSR J0437-4715, its first mass and radius are $M=1.418^{+0.037}_{-0.037}\,M_{\odot}$ and $R=11.36^{+0.95}_{-0.63}$ km, by using the 2017-2021 NICER X-ray spectral-timing data [43]. The inference results for PSR J1231-1411 exhibit a strong sensitivity to the choice of radius priors, with stable and well-converged outcomes achieved only under constrained radius priors. When the radius is restricted to agree with previous observational constraints and EoS analyses, the inferred values are $R=12.60_{-0.33}^{+0.31}$ km and $M=1.04_{-0.03}^{+0.05}M_{\odot}$ [44].
@@
The mass range for neutron stars covered by the aforementioned data spans the range 1-$2\,M_{\odot}$. Correspondingly, the relevant range for densities spans $2$-$4\,\rho_{\rm sat}$, with $\rho_{\rm sat}\simeq 0.16$ fm-3 being the nuclear saturation density. This opens an unprecedented opportunity to explore models of modern theories of dense matter subject to astrophysical constraints. For example, the impact of variations in the estimates for PSR J0030+0451 [40] on dense matter theories and CS properties have recently been assessed [10, 45, 31]. In addition, the X-ray observations of the temperature of the neutron star in the Cassiopeia A supernova remnant and modeling of its thermal evolution give insights into the possible superfluid properties of the core matter [46].
@@
Determining the maximum mass of neutron stars is crucial for understanding stellar evolution, supernovae, and binary mergers, as well as their observational signatures. This value primarily depends on the EoS at high densities, $\rho>4\rho_{\rm{sat}}$. Most observed neutron stars rotate significantly more slowly than their theoretical Keplerian break-up limit (e.g. Ref [47] and references therein). As a result, mass measurements are commonly used to establish lower bounds on the maximum mass of cold, non-rotating neutron stars $M_{\rm max}$, even though rapid rotation can support up to $\sim 20\%$ additional mass. The highest precisely and reliably measured massive mass belongs to pulsar PSR J0348+0432, with a mass estimate of $M=2.01_{-0.04}^{+0.04}\,M_{\odot}$ (at 68.3 % Cl) [48]. The heavier pulsar, PSR J0740+6620, has a mass of $M=2.08_{-0.07}^{+0.07}\,M_{\odot}$ [49, 50] and excludes all the EoS models which cannot reach this limit. At the same time the analysis of the kilonova AT2017gfo and GRB170817A [51, 52], associated with GW170817, have helped in designing theoretical models of the post-merger phase of GW170817 that placed an upper limit on the maximum mass $M_{\rm max}\sim 2.3\,M_{\odot}$ [53, 54, 55, 56, 57, 58].
@@
These mass limits have been further challenged by evidence suggesting the existence of even heavier compact objects. The “black widow” pulsar PSR J0952-0607, the second-fastest known spinning CS with a rotation frequency of 707 Hz, was detected in the Milky Way with a mass estimate of $M=2.35_{-0.17}^{+0.17}\,M_{\odot}$ (at 68.3% CI) [59], making it the most massive neutron star observed to date.
@@ I.2 Summary of nucleonic models of dense matter
The EoS for dense matter in CSs can be constructed using various approaches, including ab initio calculations [73, 74, 75, 76], density functional methods [77, 78, 79, 80], and meta-type models [81, 82, 16], which provide a more schematic representation by parametrizing nuclear saturation properties. These models can also be linked to heavy-ion collision data [83, 84].
@@
The covariant density functional (CDF) models provide a rigorous framework to address the full range of available data on nuclear systems, ranging from the atomic chart to the astrophysics of CSs, for reviews see Refs. [86, 87, 77, 79, 80]. These models were instrumental in addressing successfully such astrophysics problems as the hyperon puzzle in conjunction with two-solar mass CSs and the TD inference of the GW170817 event. The CDF models are separated into broad classes which (a) have non-linear meson contributions to the effective Lagrangian and (b) keep only linear coupling but impose density-dependence of the coupling, which captures the medium modifications of the meson-nucleon vertices. The mesons effectively considered, and their self- or cross-couplings in the former class or the explicit forms of density-dependences of the couplings in the latter class, are commonly flexible in order to address specific issues. Consequently, their predictions, are somewhat model-dependent, as different density functionals result in different dependences on density, see reviews [77, 79, 80].
@@ II Density-dependent covariant density functionals for nucleonic matter
where $\psi_{N}$ are the nucleonic Dirac fields with masses $m_{N}$, $\bm{\tau}$ is the vector of isospin Pauli matrices with $\tau_{3}$ being its third component, and $\sigma,\,\omega_{\mu}$, and $\bm{\rho}_{\mu}$ represent the mesonic fields that mediate the interaction among nucleon fields with the corresponding coupling constants $g_{\sigma},\,g_{\omega}$ and $g_{\rho}$, respectively. The remaining pieces of the Lagrangian correspond to the mesonic, and leptonic contributions, respectively [77, 80].
The particle masses adopted in the present analysis are listed in Table 1, according to the widely used DD-ME parametrization [89, 90]. Note that the mass of the $\sigma$ meson $m_{\sigma}$, which is supposed to represent the two-$\pi$-exchange contribution to the nuclear force is not known with precision and lies around 500 MeV. The properties of the infinite nuclear matter described by the present model depend only on the ratios of coupling constants and the corresponding meson masses [91, 92], and our results do not change with variation of $\sigma$ meson mass $m_{\sigma}$, as its variations are equivalently encoded into the variations of $g_{\sigma}$ as fixed $m_{\sigma}$.
Table 1: Nucleon and meson masses (in units of MeV) adopted in the CDF parametrization.
@@
The dependence of meson-nucleon couplings on the total baryon (vector) density is assumed to be
@@
which is important for a quantitative description of finite nuclei, e.g., spin-orbit splitting [94]. The rearrangement term $\Sigma_{R}$ in Eqs. (6) arises from the density dependence of meson-baryon couplings and is included to ensure thermodynamical consistency. It is seen that at zero temperature, the energy density and pressure of stellar matter are functions of baryonic density only.
It is now clear that if we fix in the Lagrangian (II) the nucleon and meson masses to be (or be close to) the ones in the vacuum then properties of infinite nuclear matter can be computed uniquely in terms of seven adjustable parameters. These are the three coupling constants at saturation density $(g_{\sigma},\,g_{\omega},\,g_{\rho})$, and four parameters $(a_{\sigma},\,d_{\sigma},\,d_{\omega},\,a_{\rho})$ that control their density dependence. We further require conditions of weak equilibrium and charge neutrality that prevail in CSs. Combined with the field equations for baryons and mesons, they allow one to determine the equilibrium composition $\rho_{N}$ and $\rho_{l}$ at a given baryon number density and, thus, determine the EoS of stellar matter.
The energy density of isospin asymmetric matter is customarily split into an isoscalar term and an isovector term:
@@
where $\chi=(\rho-\rho_{\rm{sat}})/3\rho_{\rm{sat}}$. The coefficients of the expansion are known as nuclear matter characteristics at saturation density, namely, binding energy $E_{\rm{sat}}$, incompressibility $K_{\rm{sat}}$, the skewness $Q_{\rm{sat}}$ in the isoscalar sector, and the symmetry energy $J_{\rm{sym}}$ and its slope parameter $L_{\rm{sym}}$ in the isovector sector. Given the five macroscopic characteristics together with the preassigned values of $\rho_{\rm sat}$ and $M^{\ast}_{\rm D}$, one can determine uniquely the seven adjustable parameters of the density functional defined above.